Can you have a look at the CLOSING ENTRIES at the end, where i put a red rectangle? and tell me if I answered correctly to the question? if not please provide me with the correct closing entries. ASAP, i have limited time to post this homework. Thanks

At year-end, an analysis by the city’s finance department determined the following constraints on fund balances in the General Fund. Prepare the appropriate closing/reclassification journal entry in the General Fund to reclassify amounts between Fund Balance—Unassigned and the fund balance accounts corresponding to the constraints shown below. (Note: You should consider the beginning of year balances in fund balance accounts in calculating the amounts to be reclassified. Be sure the check mark in the box for [Closing Entry] is showing before closing each individual account.)

                       Account                                                     Ending Balance

Fund Balance—Committed—General Government        $30,000

Fund Balance—Restricted—Public Safety                        36,000

            Fund Balance—Committed—Public Works                      12,700

            Fund Balance—Assigned—Culture and Recreation                   0

JULIIVINE Current accounting entity 1-General Fund Journal Ledger Accounts Entities Journal entries table Search a Add new enRevenues 2020 4-a-1 1,679,070 103 Encumbrances 2020 4-a-2 608,314 103 Encumbrances Outstanding 608,314 103 Cash 3,399,599 103  572,777 Vouchers Payable Due from State Government 4-a-5 4-a-6 4-a-6 115,000 Revenues 115,000 2020 2020 2020 2020 2020 2020 2  .. 169,465 Taxes Receivable-Delinquent Taxes Receivable-Current Allowance for Uncollectible Current Taxes 4-a-8 4-a-8 4-a-8 .
Answer #1

As per questions its clear that "We should consider the beginning of year balances in fund balance accounts in calculating the amounts to be reclassified"

So,

Fund Balance—Committed—General Government beginning balance is zero.

Fund Balance—Restricted—Public Safety beginning balance is CR 15,000

Fund Balance—Committed—Public Works beginning balance is CR 29,700

Fund Balance—Assigned—Culture and Recreation beginning balance is CR 56,800

Closing entries should be.

Fund Balance—Committed—General Government DR 30,000 "(30,000 +/- 0) = 30,000"

Fund Balance—Restricted—Public Safety DR 51,000 "(51,000-15,000(BB)) = 36,000"

Fund Balance—Committed—Public Works DR 42,400 "(42,400 -29,700(BB)) = 12,700"

Fund Balance—Assigned—Culture and Recreation DR 56,800 "(56,800-56,800(BB)) = 0"

Fund Balance - Unassigned CR 1,80,200

(Being Fund balances reinstated after considering beginning balance)

BB = beginning balance